OpenStudy (anonymous):

How would you factor x^4-81? I need to see the steps; I have the answer but not how you get it.

OpenStudy (anonymous):

\[x^4-81=(x^2)^2-9^2=(x^2-9)(x^2+9)\] \[=(x-3)(x+3)(x^2+9)\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

That was pretty nice. I would fan you, Anwar, if not for you getting so cheeky earlier!

OpenStudy (anonymous):

Newton is a fan of me.. YAAY!! :)

OpenStudy (anonymous):

When was I cheeky? I am never cheeky, if I know what that means :P